i need help to write this in slope intercept form and then graphing it.
x+3y-3=0

Answer by dabanfield(803)

i need help to write this in slope intercept form and then graphing it.
x+3y-3=0
The slope-intercept form of a line is:
y = m*x + b
where m is the slope and b is the value of y where the line intercepts the x-axis (i.e. where x = 0).
We have:
x+3y-3=0
3y = -x+3
y = (-1/3)*x +(3/3)
y = (-1/3)*x + 1
m = -1/3 and b = 1